a corrosion inhibitor is being dosed at a rate of 20 ml/min into a water distribution pipe with a water flow rate of 150 gal/min. a sample is collected and measured by uv/vis and shows a phosphate residual of 3 mg/L. if the corrosion inhibitor has a density of 1.15 g/ml, calculate its %PO4 content


1
Expert's answer
2021-03-31T08:02:56-0400

Rate of inhibitor=20ml/min

Rate of water flow= 150gal/min

Conc of phosphate= 3mg/L

Density of inhibitor= 1.15g/ml

Now,

1ml of inhibitor contains 1.15g

20ml of the inhibitor will contain 1.15/1 x 20=23g

Also,

1000ml(1L) of the inhibitor contains 3mg of PO4

20ml of the inhibitor will contain 3/1000 x 20= 0.06mg= 6x10-5g of PO4

We can now find the %w/w of PO4

%w/w= weight of PO4/weight of inhibitor x 100%

= 6x10-5g/23g x 100%=0.0003%